OpenStudy (anonymous):

Which equation is a quadratic equation? y + 3 = 5x + 10 y + 4x2 = (4x − 2)(x + 3) y − 8x = (x2 + 1)(9x − 7) y = 2x + 7x(x + 1)

OpenStudy (anonymous):

quadratic equation is equation with power of x as 2 , so here the option D

OpenStudy (anonymous):

y = 2x + 7x(x + 1)

OpenStudy (anonymous):

can u help with a few others ?

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

What is the value of the coefficient "a" when the quadratic equation y = (3x − 4)(2x − 1) is written in standard form? 6 5 −11 −4

OpenStudy (anonymous):

Can u write it in standard form by multiplying within the brackets

OpenStudy (anonymous):

y = 6x^2 -11x + 4 So, a = 6

OpenStudy (anonymous):

Identify the vertex for the graph of y = 2x2 + 8x − 3. (2, 21) (2, 17) (−2, −11) (−2, −27)

OpenStudy (anonymous):

\[y=ax^2+bx+c\] first coordinate of the vertex is \(-\frac{b}{a}\) \[y=2x^2+8x-3\] first coordinate of the vertex is \(-\frac{8}{2\times 2}=-2\)

OpenStudy (anonymous):

sorry i meant first coordinate of the vertex is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

@satellite73 so 2, -11 ?